May 30, 2019 Calorimeter

From GlueXWiki
Jump to: navigation, search

Meeting Time: 11:00 a.m.

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone in CC L207 is 757-269-7084 (usual room)
  • JLab Phone: in CC F326 is 757-269-6460
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book

Goals for Calorimetry Group

  1. Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
  2. Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at east 5 MeV.
  3. Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.

Action Items

Short term items:

  1. Complete Calorimeter work packages

Long term items:

  1. From Mike's work: -cos(phi) type dependence in the data and a phi dependence at the downstream end of the bcal. See Sep 21, 2017 Calorimeter Meeting.
  2. FCAL: It has been noticed that the pi0 mass shifts when the LED pulser runs at high rate. This is still not understood. See Log Entry 3595928 and references. Several runs have been taken in raw mode. These runs exhibit the shift in pi0 mass with LED running at 1 kHz. There is a need to analyze these to try to determine the cause:
    1. 50634 - FCAL Green LED: 1kHz, 29 V, Livetime is 9%, 1.7 M events
    2. 50635 - FCAL Violet LED: 1kHz, 22 V, Livetime is 9%, 1.4 M events
    3. 50636 - FCAL Blue LED: 1kHz, 15 V, Livetime is 9%, 1.5 M events

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Run Updates
    1. FCAL
    2. BCAL
  4. Calibrations
  5. Monitoring
  6. Simulations
  7. BCAL Deep Learning
  8. NIM articles
  9. Any other business


Attending: Mark, Sean Colin (JLab), Will (CMU), Matt (IU), Christina, George (Athens), Tegan, Karthik, Zisis (Regina)

  1. Announcements: none
  2. Action Items
    1. Work packages: (item number below refer to the table on that Wiki page)
      1. Mark reported that the first iteration of this item has been completed, employing the simplest cuts. More sophisticated cuts will be investigated in the future. Sean sent an email last week: he implemented the changes but needs someone to review them before they are checked in.
      2. Karthik is working on the report and it should be available soon. The BCAL response near its ends is not completely understood.
      3. Colin reported that item 3 has been completed.
      4. Matt reported that Jon is working on this and should be able to report progress in two weeks. The next thing to tackle is the implementation of fiducial cuts; Sean has this in a branch.
      5. Karthik will start looking at this after his report on Item 2 is done.
      6. Mark suggested that this item needs the previous to be completed and then can act as a stress test on the method. Preliminary studies indicated that we can access some parts of phase space.
      7. Nothing new on this item.
      8. Matt said that work is ongoing on the MC although not in connection to energy resolution. The work is focused on tuning the neural net and nonlinearity. Data tables will be added to CCDB to tune the low energy response of the FCAL (Rebecca). Sean also reported that Lawrence (FSU student) will work on the resolution comparison after the HUGS school.
      9. Nothing to report.
      10. Tolga has started working on the timing (neutrons vs photons) and has an iniitlal list of things to purse, but is now in HUGS. He should be able to report in 4-6 weeks.
    2. Mike's cos(theta): these investigations are included in the work packages.
    3. FCAL pulser: nothing to report.
  3. Run Updates
    1. FCAL: a test of the FADC firmware is planned. Sean will look into it.
    2. BCAL: Mark mentioned the campaign getting underway to extract the BCAL pixel/ADC count. Two Regina students will be at JLab for three weeks to work on this. Fernando's lab will be used. There is a working DAQ, dark box and laser. Mark will see if there is material for the students to read before arriving at the lab. They have completed most training, with the exception of laser, Rad Worker I and medical exam, which they will do at the lab.
  4. Calibrations
    1. FCAL: 4ns offset calibrations for fall 2018 are completed, and spring 2019 is close to completion as well.
    2. BCAL: Karthik will start spring 2019 pi0 calibrations on skims, as a first pass, before lowe level calibrations are completed. Mark noted that there are two distinct data sets, one with the solenoid off. Sean suggested to use the existing (fall 2018) calibrations for spring 2019 and see if there are significant differences, as the BCAL was used mostly as a veto in Primex and we may not need to put in a lot of effort for the spring calibration.
  5. Monitoring:
    1. BCAL: Ahmed uploaded a new version of his LED-Expert monitoring document to DocDB (doc no 3975). This is a dynamic document and will be updated as necessary. The plan is to have this system fully operational for fall 2019 running.
    2. FCAL: An IU summer student is looking at the LED info with the aim to look for radiation damage and assess the stability of the inner rings (for Primex). Mark will send the location of a Wiki page to Matt, with information on the LED colors and voltages; these are currently tied to the wall clock and have 6 pulsing options, each for 10 min within an hour.
  6. Simulations: Mark and Richard are working on understanding a 2% difference in the BCAL overall-factor calibration in G4 when compared to G3.
  7. BCAL Deep Learning: Zisis provided an overview of the objectives for particle ID and clusterizer, as well as comparing different neural nets (XGBoost, TensorFlow). Meetings are weekly, at 1 pm JLab time. Tegan completed the resurrection of the BDT/TVMA work that he and Justin had done in 2015. This required considerable work since the GlueX framework changed significantly since then. The BDT classifies showers as good (if matched to thrown). Everything else is treated as noise. Results look ok and comparable to 2015 in terms of signal to noise separation.
  8. NIM articles
  9. Any other business: none.